OpenStudy (anonymous):

During a family trip, you share the driving with your dad. At most, you are allowed to drive for three hours. While driving, your maximum speed is 55 miles per hour. a) Write a system of inequalities describing the possible numbers of hours t and distance d you may have to drive. b) Is it possible for you to have driven 160 miles?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@dan815

jimthompson5910 (jim_thompson5910):

hopefully you recall that distance = speed * time agreed?

OpenStudy (anonymous):

I agree

jimthompson5910 (jim_thompson5910):

let d = distance r = rate or speed t = time

OpenStudy (anonymous):

Ok

jimthompson5910 (jim_thompson5910):

"While driving, your maximum speed is 55 miles per hour" so `r <= 55`. The value of r must be smaller than or equal to 55 mph

OpenStudy (anonymous):

Ok

jimthompson5910 (jim_thompson5910):

In addition, `r >= 0` because we can't have a negative speed

OpenStudy (anonymous):

Ok

jimthompson5910 (jim_thompson5910):

` At most, you are allowed to drive for three hours` telling us that `t <= 3`. For similar reasons as above, `t >= 0` as well

OpenStudy (anonymous):

Ok

jimthompson5910 (jim_thompson5910):

so we have this so far r <= 55 r >= 0 t <= 3 t >= 0 let's see. what else is there

OpenStudy (anonymous):

b) Is it possible for you to have driven 160 miles?

jimthompson5910 (jim_thompson5910):

if your max speed is 55 mph and the max amount of time is 3 hrs, then d = r*t d = 55*3 d = 165 so the max distance possible is 165 miles. It's definitely possible to drive 160 miles

OpenStudy (anonymous):

Thank you so much!

jimthompson5910 (jim_thompson5910):

np